# Quandles as pre-Lie skew braces, set-theoretic Hopf algebras &\&& universal ${\cal R}$-matrices

### v1

@@ 3.2. Generalised affine shelves and Yang-Baxter algebras
    [27, Definition 1] $f$ is said to be a metahomomorphism of $\left(X,+\right)$ if
@@
    if $f$ is a metahomomorphism of $(X,+)$, then $\hat{f}$ is metahomomorphism of $(X,+)$ such that $f(0)=0$, namely, $f$ is a unitary metahomomorphism. More generally, any metahomomorphism can be obtained from a unitary one (cf. [27, Theorem 4]).
@@
We are now ready to introduce generalisations of affine quandles. We include the first example as it was inspired by the affine quandle. Compare with [29, (1.1)] and [8, Theorem 5.9].
@@
is a spindle, and if $f$ is bijective it is a quandle. More generally, if $f:X\to X$ is a bijective map such that $f(0)=0$, it is easy to check that $\left(X,\triangleright_{t}\right)$ is a quandle if and only if $f\left(a-f\left(b\right)+b\right)=f\left(a\right)-f^{2}\left(b\right)+f\left(b\right)$ holds, for all $a,b\in X$. Note that, in the case of $f\in\operatorname{Aut}\left(X,+\right)$, the structure $\left(X,\triangleright_{t}\right)$ coincides with (right) quandles provided in [30, p. 42].
@@
is a solution of the braid equation if and only $f$ is a metahomomorphism of $\left(X,+\right)$, see [27, Theorem 1]. Analogously, one can check that the map $s:X\times X\to X\times X$ defined by
@@
[…]
@@ 2. Right pre-Lie condition
We note that the distributive condition in Definition 3.23 appears naturally in skew braces, also the right pre-Lie condition is satisfied in Definition 3.23, hence the name right pre-Lie skew braces for the structure described in Definition 3.23.
@@
Recall that for the pre-Lie skew braces above: $a\bullet b=b\bullet(b\triangleright a),$ where the normal-$\triangleright$ action is given in 3.14. Because of the existence of the left inverse it follows:
@@
It is a simple check that $(P,[-,-,-],\{-,-,-\})$ satisfies conditions of [6, Definition 3.1]
@@
Immediately follows by Lemma 3.30 and [6, Proposition 3.6]. ∎
@@
4. Deformed braided algebras &\&& racks
Motivated by the definition of braided groups and braidings in [35] as well as the relevant work presented in [25] and the generic definition of deformed braided algebras and deformed braidings that contain extra fixed parameters [18, 19] we provide a relevant definition associated to racks.
First, let us fix some notation. In this section, we will fix the following arbitrary invertible maps for any set $X$
@@ 4. Deformed braided algebras &\&& racks
[…]
@@
[…]
@@
We note that given the existence of an invertible, non-degenerate, set-theoretic solution of the braid equation, together with the group $(X,\circ),$ the construction of a skew brace consistently follows (see details of such a construction in [18, Section 2.2].)
Motivated by recent observations on deformed solutions and connections to Yang-Baxter algebras from section 2, we introduce the following notion of a generalised braiding operator designed for Yang-Baxter algebras. Observe that, we put a condition on the form of the braiding, but this condition naturally arises from the construction of a quandle/shelf solution.
@@
The map $r$ is called an $S$-deformed braiding operator (and the algebra is called $S$-deformed braided algebra) if for all $x,y,w\in X:$
@@
[…]
@@
[…]
@@
Let $(X,\bullet)$ be an $S$-deformed braided algebra and the map $r:X\times X\to X\times X,$ such that $r(x,y)=(y,\tau_{y}(x))$ be an $S$-deformed braiding operator, then $r$ satisfies the braid equation.
@@
Let $(X,\bullet)$ be a $S$-deformed braided algebra and the map $r:X\times X\to X\times X,$ such that $r(x,y)=(\sigma_{x}(y),x),$ be a $S$-deformed braiding operator, then $r$ satisfies the braid equation.
@@
Let $(X,+{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@rgb@stroke{0}{1}{0}\pgfsys@color@rgb@fill{0}{1}{0},}\circ)$ be a skew brace, $\sigma_{a}(b)=-a+a\circ b,$ $\tau_{b}(a)=\sigma_{a}(b)^{-1}\circ a\circ b,$ and $r(a,b)=(\sigma_{a}(b),\ \tau_{b}(a)).$ In this case $\sigma_{a}(\sigma_{b}(c))=\sigma_{a\circ b}(c),$ similarly $\tau_{c}(\tau_{b}(a))=\tau_{b\circ c}(a),$ i.e. $\sigma_{a}=f_{a}=\hat{f}_{a}.$ Thus $r$ is a braiding and $(X,\circ)$ is a braided group.
@@
Let $(X,+)$ be a group, $(X,\triangleright_{s})$ the quandle in 3.14, and $\tau_{b}(a):=b\triangleright_{s}a=-f(b)+f(a)+b$, for all $a,b\in X$, with $f$ a bijective heap endomorphism of $(X,+)$. We also consider the binary operation on $X$ in Lemma 3.20 such that $a\bullet b=f(a)+b$, for all $a,b\in X$. In this case $\tau_{c}(\tau_{b}(a))=\tau_{b\bullet c}(a\bullet e)$, i.e. $\hat{g}_{b}(a)=\tau_{b}(a\bullet e),$ where $e=f^{-1}(0)$ is the left neutral element in $(X,\bullet).$ Then, $r(a,b)=(b,\ \tau_{b}(a))$ is an $S$-deformed braiding and $(X,\bullet)$ is an $S$-deformed braided group.
@@ 5. Quasi triangular (quasi)-Hopf algebras
We discuss in this section the Yang-Baxter algebras associated to quandle type and set-theoretic solutions as quasitriangular (quasi)-Hopf algebras [21, 22]. The quasi Hopf algebra for set-theoretic solutions has been discussed in [16, 17, 18] (see also [23] and [2] in connection with pointed Hopf algebras), whereas a discussion of bialgebras associated to racks is presented in [15] (see also [31, 3] on Hopf algebras in connection to braces [36][38], [28]).
We start our analysis with the quandle Hopf algebra and the construction of the associated universal ${\cal R}$-matrix. We then extend the algebra and construct the set-theoretic Hopf algebra and via a suitable admissible Drinfel’d twist we construct the corresponding universal ${\cal R}-$matrix.
5.1. The quandle Hopf algebra
We first define the quandle algebra (see also [2, 15], where part of the defined algebra below is studied).
@@ 5.1. The quandle Hopf algebra
(The quandle algebra.) Let $(X,\triangleright)$ and $(X,\bullet)$ be left quasigroups, such that, for all $a,b\in X,$ $a\bullet b=b\bullet(b\triangleright a)$. We say that the unital, associative algebra ${\cal A},$ over a field $k$ generated by indeterminates ${1A},\ q_{a},\ q^{-1}_{a},\ h_{a}\in{\cal A}$ ($h_{a}=h_{b}\Leftrightarrow a=b$) and relations, $a,b\in X:$
$q_{a}q_{a}^{-1}=1_{\cal A},\quad q_{a}q_{b}=q_{b}q_{b\triangleright a},\quad h_{a}h_{b}=h_{b}h_{a}=\delta_{a,b}h_{a}^{2},\quad q_{b}h_{b\triangleright a}=h_{a}q_{b},$ (5.1)
is a quandle algebra.
Lemma 5.2.
Let $c=\sum_{a\in X}h_{a},$ then $c$ is a central element of the quandle algebra ${\cal A}$.
@@
The proof is straightforward by means of the definition of the algebra ${\cal A}.$
We consider henceforth, without loss of generality, $c=1_{\cal A},$ from the latter it also immediately follows that $\sum_{a\in X}h_{a}^{2}=1_{{\cal A}}.$
Proposition 5.3.
Let ${\cal A}$ be the quandle algebra (Definiton 5.1,) then $c\triangleright(b\triangleright a)=(c\triangleright b)\triangleright(c\triangleright a)$ and $a\triangleright a=a,$ i.e. $(X,\triangleright)$ is a quandle
Proof.
We compute $h_{a}q_{b}q_{c}$ using the associativity of the algebra:
@@
Moreover, we recall $a\bullet b=b\bullet(b\triangleright a),$ then for $a=b$ and by recalling bijectivity and hence left cancellativity in $(X,\bullet)$ we conclude that $a\triangleright a=a,$ i.e. $(X,\triangleright)$ is a quandle. ∎
@@
Let ${\cal A}$ be the quandle algebra, then $q:X\to{\cal A},$ $a\mapsto q_{a}$ is called:
@@
Having defined the quandle algebra we are now in the position to identify the associated universal ${\cal R}$-matrix (solution of the Yang-Baxter equation).
@@
Let ${\cal A}$ be the quandle algebra and ${\cal R}\in{\cal A}\otimes{\cal A}$ be an invertible element, such that ${\cal R}=\sum_{a}h_{a}\otimes q_{a}.$ Then ${\cal R}$ satisfies the Yang-Baxter equation
@@
$\mbox{LHS}:\quad\sum_{a,b,c\in X}h_{a}h_{b}\otimes q_{a}h_{c}\otimes q_{b}q_{c}=\sum_{a,b,c\in X}h_{a}^{2}\otimes q_{a}h_{c}\otimes q_{a}q_{c}=\sum_{a,b,c\in X}h^{2}_{a}\otimes q_{a}h_{a\triangleright c}\otimes q_{a}q_{a\triangleright c}$
$\mbox{RHS}:\quad\sum_{a,b,c\in X}h_{b}h_{a}\otimes h_{c}q_{a}\otimes q_{c}q_{b}=\sum_{a,b,c\in X}h^{2}_{a}\otimes q_{a}h_{a\triangleright c}\otimes q_{c}q_{a},$
and due to the basic relation $q_{a}q_{b}=q_{b}q_{b\triangleright a},$ we show that LHS$=$RHS, and this concludes our proof. ∎
@@
We require ${\cal R}$ to be invertible, but from Lemma 5.2, $\sum_{a\in X}h^{2}_{a}=1_{\cal A},$ for all $a\in X,$ hence ${\cal R}^{-1}=\sum_{a\in X}h_{a}\otimes q_{a}^{-1}.$
@@
[…]
@@
Let ${\cal P}=\sum_{a,b\in X}e_{a,b}\otimes e_{b,a}$ be the permutation (flip) operator, then the solution of the braid equation is the familiar quandle solution, $r={\cal P}R=\sum_{a,b}e_{a,b}\otimes e_{b,b\triangleright a}.$ We note that for $R$ to be invertible, $a\triangleright:X\to X$ should be a bijection, which is compatible with Definition 5.1, then $R^{-1}=\sum_{a,b\in X}e_{b,b}\otimes e_{b\triangleright a,a}.$
@@
$(\pi\otimes\mbox{id}){\cal R}:=L=\sum_{a\in X}e_{a,a}\otimes q_{a},$ $\leavevmode\nobreak\ (\mbox{id}\otimes\pi){\cal R}:=\hat{L}=\sum_{a,b\in X}h_{b}\otimes e_{a,b\triangleright a},$ and
$(\pi\otimes\pi){\cal R}:=R=\sum_{a,b\in X}e_{b,b}\otimes e_{a,b\triangleright a},$ consistent algebraic relations ensue:
@@
which lead to the quandle algebra given in Definition (5.1) and provide a consistency check on the algebraic relations (5.1). This is the Faddeev-Reshetikhin-Takhtajan (FRT) construction [24].
@@
[…]
@@
[…]
@@
[…]
@@
    If $(X,\bullet,e)$ is a group with identity $e$ and inverse denoted by $a^{*},$ for every $a\in X,$ then $b\triangleright(a\triangleright x)=(a\bullet b)\triangleright x,$ for every $x\in X,$ and consequently $\rho_{a}\rho_{b}=\rho_{a\bullet b},$ i.e. this is a strong group homomorphism. In this case $\rho_{a^{*}}=\rho^{-1}_{a}$ and $\rho_{e}=\pi(1_{\cal A})=id_{V}.$ Notice also that $\rho_{a}=\rho_{b}\Rightarrow x\bullet(a\bullet b^{*})=(a\bullet b^{*})\bullet x,$ for every $x\in X,$ i.e. $a\bullet b^{*}$ is a central element in the group $(X,\bullet).$
@@
    If $(X,\bullet_{s})$ is the pre-Lie skew brace studied in subsection 3.3, with $a\triangleright_{s}b=-f(a)+f(b)+a,$ $\leavevmode\nobreak\ a\bullet_{s}b=f(a)+b,$ then $b\triangleright_{s}(a\triangleright_{s}x)=(a\bullet_{s}b)\triangleright_{s}(x\bullet_{s}e),$ for every $x\in X$ (recall $e=f^{-1}(0)$) and consequently $\rho_{a}\rho_{b}=F_{a\bullet_{s}b}:=\sum_{x\in X}e_{x,(a\bullet_{s}b)\triangleright_{s}(x\bullet_{s}e)},$ i.e. this is a weak algebra homomorphism. In this case $\rho_{a}^{T}=\rho^{-1}_{a}$ (${}^{T}$ denotes transposition) and $\sum_{a\in x}e_{a,a,}=\pi(1_{\cal A})=id_{V}.$ Notice also that if $\rho_{a}=\rho_{b}\Rightarrow a-b=-f(x)+f(a)-f(b)+f(x),$ for every $x\in X.$

### v3

@@ 3.2. Generalized affine shelves and Yang-Baxter algebras
    [37, Definition 1] $f$ is said to be a metahomomorphism of $\left(X,+\right)$ if
@@
    if $f$ is a metahomomorphism of $(X,+)$, then $\hat{f}$ is metahomomorphism of $(X,+)$ such that $f(0)=0$, namely, $f$ is a unitary metahomomorphism. More generally, any metahomomorphism can be obtained from a unitary one (cf. [37, Theorem 4]).
@@
We are now ready to introduce generalisations of affine quandles. We include the first example as it was inspired by the affine quandle. Compare with [40, (1.1)] and [10, Theorem 5.9].
@@
is a spindle, and if $f$ is bijective it is a quandle. More generally, if $f:X\to X$ is a bijective map such that $f(0)=0$, it is easy to check that $\left(X,\triangleright_{t}\right)$ is a quandle if and only if $f\left(a-f\left(b\right)+b\right)=f\left(a\right)-f^{2}\left(b\right)+f\left(b\right)$ holds, for all $a,b\in X$. Note that, in the case of $f\in\operatorname{Aut}\left(X,+\right)$, the structure $\left(X,\triangleright_{t}\right)$ coincides with (right) quandles provided in [45, p. 42].
@@
is a solution of the braid equation if and only $f$ is a metahomomorphism of $\left(X,+\right)$, see [37, Theorem 1]. Analogously, one can check that the map $s:X\times X\to X\times X$ defined by
@@
$a\triangleright_{r}b=a+f\left(b\right)-f\left(a\right)\qquad\text{and}\qquad a\triangleright_{s}b=-f\left(a\right)+f\left(b\right)+a,$
@@ 2. Right pre-Lie condition
We note that the distributive condition in Definition 3.23 appears naturally in skew braces, also the right pre-Lie condition is satisfied in Definition 3.23, hence the name right pre-Lie skew braces is well justified for the structure described in Definition 3.23.
@@
Recall that for the pre-Lie skew braces above: $a\bullet b=b\bullet(b\triangleright a),$ where the $\triangleright$ action is given in 3.14. Because of the existence of the left inverse it follows:
@@
It is a simple check that $(P,[-,-,-],\{-,-,-\})$ satisfies conditions of [11, Definition 3.1]
@@
Immediately follows by Lemma 3.30 and [11, Proposition 3.6]. ∎
@@
4. Deformed braided algebras & magmas
Motivated by the definition of braided groups and braidings in [49] as well as the relevant work presented in [33] and the generic definition of deformed braided algebras and deformed braidings that contain extra fixed parameters [23, 24] we provide a relevant definition associated to magmas.
It is useful for the following definition to fix the arbitrary invertible maps for any set $X:$
@@ 4. Deformed braided algebras & magmas
    $\xi(m\times\mbox{id}_{X})(x,y,w)=(\mbox{id}_{X}\times m)(r\times\mbox{id}_{X})(\mbox{id}_{X}\times r)(x,y,w).$
@@
    $\zeta(\mbox{id}_{X}\times m)(x,y,w)=(m\times\mbox{id}_{X})(\mbox{id}_{X}\times r)(r\times\mbox{id}_{X})(x,y,w).$
@@
We note that given the existence of an invertible, non-degenerate, set-theoretic solution of the braid equation, together with the group $(X,\circ),$ the construction of a skew brace consistently follows (see details of such a construction in [23, Section 2.2].)
Motivated by recent observations on deformed solutions and connections to Yang-Baxter algebras from section 2, we introduce the following notion of a generalized braiding operator designed for Yang-Baxter algebras. Observe that, we put a condition on the form of the braiding, but this condition naturally arises from the construction of a quandle/shelf solution.
@@
The map $r$ is called an $S$-deformed braiding operator (and the magma is called $S$-deformed braided magma) if for all $x,y,w\in X:$
@@
    $\xi(m\times\mbox{id}_{X})(x,y,w)=(\mbox{id}_{X}\times m)(r\times\mbox{id}_{X})(\mbox{id}_{X}\times r)(x,y,w).$
@@
    $\zeta(\mbox{id}_{X}\times m)(x,y,w)=(m\times\mbox{id}_{X})(\mbox{id}_{X}\times r)(r\times\mbox{id}_{X})(x,y,w).$
@@
Let $(X,\bullet)$ be an $S$-deformed braided magma and the map $r:X\times X\to X\times X,$ such that $r(x,y)=(y,\tau_{y}(x))$ be an $S$-deformed braiding operator, then $r$ satisfies the braid equation.
@@
Let $(X,\bullet)$ be an $S$-deformed braided magma and the map $r:X\times X\to X\times X,$ such that $r(x,y)=(\sigma_{x}(y),x),$ be an $S$-deformed braiding operator, then $r$ satisfies the braid equation.
@@
Let $(X,+,\circ)$ be a skew brace, $\sigma_{a}(b)=-a+a\circ b,$ $\tau_{b}(a)=\sigma_{a}(b)^{-1}\circ a\circ b,$ and $r(a,b)=(\sigma_{a}(b),\ \tau_{b}(a)).$ In this case $\sigma_{a}(\sigma_{b}(c))=\sigma_{a\circ b}(c),$ similarly $\tau_{c}(\tau_{b}(a))=\tau_{b\circ c}(a),$ i.e. $\sigma_{a}=f_{a}=\hat{f}_{a}.$ Thus $r$ is a braiding and $(X,\circ)$ is a braided group.
@@
Let $(X,+)$ be a group, $(X,\triangleright_{s})$ the quandle in 3.14, and $\tau_{b}(a):=b\triangleright_{s}a=-f(b)+f(a)+b$, for all $a,b\in X$, with $f$ a bijective heap endomorphism of $(X,+)$. We also consider the binary operation on $X$ in Lemma 3.20 such that $a\bullet b=f(a)+b$, for all $a,b\in X$. In this case $\tau_{c}(\tau_{b}(a))=\tau_{b\bullet c}(a\bullet e)$, i.e. $\hat{g}_{b}(a)=\tau_{b}(a\bullet e),$ where $e=f^{-1}(0)$ is the left neutral element in $(X,\bullet).$ Then, $r(a,b)=(b,\ \tau_{b}(a))$ is an $S$-deformed braiding and $(X,\bullet)$ is an $S$-deformed braided magma.
@@ 5. Quasi triangular (quasi)-Hopf algebras
We discuss in this section the Yang-Baxter algebras associated to rack type and set-theoretic solutions as quasitriangular (quasi)-Hopf algebras [27, 28]. The quasi Hopf algebra for set-theoretic solutions has been partly discussed in [21, 22, 23] (see also [29] and [3] in connection with pointed Hopf algebras), whereas a discussion of bialgebras associated to racks is presented in [19] (see also [47, 4] on Hopf algebras in connection to braces [54][56], [38]).
We start our analysis with the rack and quandle algebras and the construction of the associated universal ${\cal R}$-matrix. We then extend the algebra to the decorated rack algebra and via a suitable admissible Drinfel’d twist we construct the corresponding universal ${\cal R}-$matrix. We note that in this section we consider finite or infinite countable sets.
5.1. The rack & quandle algebras
We first define the rack and quandle algebras (see also [3, 19], where part of the defined algebra below is studied).
@@ 5.1. The rack & quandle algebras
(The rack algebra) Let $(X,\triangleright)$ be a finite magma, or such that $a\triangleright$ is surjective, for every $a\in X$. We say that the unital, associative algebra ${\cal A},$ over a field $k$ generated by indeterminates ${1A}$ (the unit element), $\leavevmode\nobreak\ q_{a},\ q^{-1}_{a},\ h_{a}\in{\cal A}$ ($h_{a}=h_{b}\Leftrightarrow a=b$) and relations, $a,b\in X:$
$q_{a}^{-1}q_{a}=q_{a}q_{a}^{-1}=1_{\cal A},\quad q_{a}q_{b}=q_{b}q_{b\triangleright a},\quad h_{a}h_{b}=\delta_{a,b}h_{a}^{2},\quad q_{b}h_{b\triangleright a}=h_{a}q_{b},$ (5.1)
is a rack algebra.
Definition 5.2.
(The quandle algebra) A rack algebra is called a quandle algebra if there is a left quasigroup $(X,\bullet)$ such that for all $a,b\in X$ $a\bullet b=b\bullet(b\triangleright a).$
The following proposition fully justifies the appellation rack and quandle algebras.
Proposition 5.3.
Let ${\cal A}$ be the rack algebra, then $c\triangleright(b\triangleright a)=(c\triangleright b)\triangleright(c\triangleright a),$ and $a\triangleright$ is bijective for all $a\in X,$ i.e. $(X,\triangleright)$ is a rack. If ${\cal A}$ is the quandle algebra, then in addition $a\triangleright a=a,$ i.e. $(X,\triangleright)$ is a quandle.
@@
We compute $h_{a}q_{b}q_{c}$ using the associativity of the rack algebra:
@@
We assume $c\triangleright a=c\triangleright b$, then $q_{c}h_{c\triangleright a}=q_{c}h_{c\triangleright b}$, by the fourth relation in 5.1, we get $h_{a}q_{c}=h_{b}q_{c}$ and by the invertibility of $q_{c}$, $h_{a}=h_{b},$ hence $a=b$, i.e. $a\triangleright$ is bijective and thus $(X,\triangleright)$ is a rack.
Moreover, we recall for the quandle algebra $a\bullet b=b\bullet(b\triangleright a),$ then for $a=b$ and by recalling bijectivity and hence left cancellativity in $(X,\bullet)$ we conclude that $a\triangleright a=a,$ i.e. $(X,\triangleright)$ is a quandle. ∎
Note: In the case of a quandle algebra $\mathcal{A}$ (without assuming $X$ to be finite), if $\left(X,\bullet\right)$ is a quasigroup, $c\triangleright$ is surjective. Indeed, if $b\in X$, and $r_{c}$ and $l_{c}$ denotes right and left multiplication by $c$, respectively, with respect to the binary operation $\bullet$, then there exists $x:=r^{-1}_{c}l_{c}\left(b\right)$ such that $c\triangleright x=l^{-1}_{c}\left(x\bullet c\right)=b$. In this case, we also obtain injectivity without using relations (5.1). Indeed, if $x,y\in X$ are such that $c\triangleright x=c\triangleright y$, then $l^{-1}_{c}r_{c}\left(x\right)=l^{-1}_{c}r_{c}\left(y\right)$, hence $x=y$. Finally, as in the last part of the proof of 5.3, we obtain that $\left(X,\triangleright\right)$ is a quandle.
Lemma 5.4.
Let $c=\sum_{a\in X}h_{a},$ then $c$ is a central element of the rack algebra ${\cal A}$. Also, $h_{a}^{2}=h_{a},$ for all $a\in X.$
Proof.
The proof is direct by means of the definition of the algebra ${\cal A}$ and Proposition 5.3. Without loss of generality let $c=1_{\cal A},$ then it also follows that $h_{a}^{2}=h_{a},$ for all $a\in X.$
@@
Let ${\cal A}$ be a quandle algebra, then $q:X\to{\cal A},$ $a\mapsto q_{a}$ is called:
@@
Having defined the rack algebra we are now in the position to identify the associated universal ${\cal R}$-matrix (solution of the Yang-Baxter equation).
@@
Let ${\cal A}$ be the rack algebra and ${\cal R}\in{\cal A}\otimes{\cal A}$ be an invertible element, such that ${\cal R}=\sum_{a\in X}h_{a}\otimes q_{a}.$ Then ${\cal R}$ satisfies the Yang-Baxter equation
@@
$\mbox{LHS}:\quad\sum_{a,b,c\in X}h_{a}h_{b}\otimes q_{a}h_{c}\otimes q_{b}q_{c}=\sum_{a,b,c\in X}h_{a}\otimes q_{a}h_{c}\otimes q_{a}q_{c}=\sum_{a,b,c\in X}h_{a}\otimes q_{a}h_{a\triangleright c}\otimes q_{a}q_{a\triangleright c}$
$\mbox{RHS}:\quad\sum_{a,b,c\in X}h_{b}h_{a}\otimes h_{c}q_{a}\otimes q_{c}q_{b}=\sum_{a,b,c\in X}h_{a}\otimes q_{a}h_{a\triangleright c}\otimes q_{c}q_{a},$
where we have used that $a\triangleright$ is bijective. Then due to the basic relation $q_{a}q_{b}=q_{b}q_{b\triangleright a},$ we show that LHS$=$RHS, and this concludes our proof. ∎
@@
The universal ${\cal R}-$matrix is obviously invertible, indeed from Lemma 5.4, $\sum_{a\in X}h_{a}=1_{\cal A},$ hence ${\cal R}^{-1}=\sum_{a\in X}h_{a}\otimes q_{a}^{-1}.$
@@
Fundamental representation: Let ${\cal A}$ be the rack algebra and $\rho:{\cal A}\to\mbox{End}(V)$ be the map defined by
@@
Let ${\cal P}=\sum_{a,b\in X}e_{a,b}\otimes e_{b,a}$ be the permutation (flip) operator, then the solution of the braid equation is the familiar rack solution, $r={\cal P}R=\sum_{a,b}e_{a,b}\otimes e_{b,b\triangleright a}.$ We note that $R$ is invertible, because $a\triangleright:X\to X$ is a bijection from Definition 5.1, then $R^{-1}=\sum_{a,b\in X}e_{b,b}\otimes e_{b\triangleright a,a}.$
@@
$(\rho\otimes\mbox{id}){\cal R}:=L=\sum_{a\in X}e_{a,a}\otimes q_{a},$ $\leavevmode\nobreak\ (\mbox{id}\otimes\rho){\cal R}:=\hat{L}=\sum_{a,b\in X}h_{b}\otimes e_{a,b\triangleright a},$ and
$(\rho\otimes\rho){\cal R}:=R=\sum_{a,b\in X}e_{b,b}\otimes e_{a,b\triangleright a},$ consistent algebraic relations ensue:
@@
which lead to the rack algebra given in Definition (5.1) and provide a consistency check on the algebraic relations (5.1). This is the Faddeev-Reshetikhin-Takhtajan (FRT) construction [32].
@@
In general, let ${\cal R}=\sum_{i}f_{i}\otimes g_{i}\in{\cal C}\otimes{\cal C}$ (${\cal C}$ is the associated unital universal algebra) be a universal ${\cal R}$-matrix. Let also, $\rho:{\cal C}\to\mbox{End}(V),$ such that $(\rho\otimes\mbox{id}){\cal R}=:L$ and $(\rho\otimes\rho){\cal R}=:R,$ then the Yang-Baxter equation reduces to (5.5). Expression (5.5) is equivalent to (3.1), if a vector $v=\sum_{a}f_{a}\hat{e}_{a}\in V,$ $f_{a}\in\mathbb{C}$ exists, such that $Rv\otimes v=v\otimes v.$ If the latter holds, and if we set $Lv=:\hat{\mathrm{q}},$ we conclude
@@
where we recall the notation: $v_{1}=v\otimes\mbox{id}\otimes 1_{\cal C},$ $v_{2}=\mbox{id}\otimes v\otimes 1_{\cal C}.$ And if $r:={\cal P}R,$ where ${\cal P}$ is the permutation operator, then (5.8) reads as, $\leavevmode\nobreak\ r_{12}\hat{\mathrm{q}}_{1}\hat{\mathrm{q}}_{2}=\hat{\mathrm{q}}_{1}\hat{\mathrm{q}}_{2}.$
@@
Let ${\cal A}$ be the quandle algebra and $\rho:{\cal A}\to\mbox{End}(V)$ the fundamental representation $\pi:{\cal A}\to\mbox{End}(V),$ $q_{a}\mapsto{\mathrm{q}}_{a}:=\sum_{b\in X}e_{b,a\triangleright b},$ $h_{a}\mapsto e_{a,a}:$
@@
    If $(X,\bullet,e)$ is a group with identity $e$ and inverse denoted by $a^{*},$ for every $a\in X,$ then $b\triangleright(a\triangleright x)=(a\bullet b)\triangleright x,$ for every $x\in X,$ and consequently $\rho_{a}\rho_{b}=\rho_{a\bullet b},$ i.e. this is a strong group homomorphism. In this case ${\mathrm{q}}_{a^{*}}={\mathrm{q}}^{-1}_{a}$ and ${\mathrm{q}}_{e}={\rho}(1_{\cal A})=id_{V}.$ Notice also that ${\mathrm{q}}_{a}={\mathrm{q}}_{b}\Rightarrow x\bullet(a\bullet b^{*})=(a\bullet b^{*})\bullet x,$ for every $x\in X,$ i.e. $a\bullet b^{*}$ is a central element in the group $(X,\bullet).$
@@
    If $(X,\bullet_{s})$ is the pre-Lie skew brace studied in subsection 3.3, with $a\triangleright_{s}b=-f(a)+f(b)+a,$ $\leavevmode\nobreak\ a\bullet_{s}b=f(a)+b,$ then $b\triangleright_{s}(a\triangleright_{s}x)=(a\bullet_{s}b)\triangleright_{s}(x\bullet_{s}e),$ for every $x\in X$ (recall $e=f^{-1}(0)$) and consequently ${\mathrm{q}}_{a}{\mathrm{q}}_{b}=F_{a\bullet_{s}b}:=\sum_{x\in X}e_{x,(a\bullet_{s}b)\triangleright_{s}(x\bullet_{s}e)},$ i.e. this is a weak algebra homomorphism. In this case ${\mathrm{q}}_{a}^{T}={\mathrm{q}}^{-1}_{a}$ (T denotes transposition) and $\sum_{a\in x}e_{a,a,}=\rho(1_{\cal A})=id_{V}.$ Notice also that if ${\mathrm{q}}_{a}={\mathrm{q}}_{b}\Rightarrow a-b=-f(x)+f(a)-f(b)+f(x),$ for every $x\in X.$